A rifle used for target shooting has a small telescope so that the shooter can accurately aim at his target. Suppose the telescope has a diameter of 1.5 cm and the shooter is looking at two small birds that are 200 m away. Assuming a red light of wavelength 600 nm shining on the birds, calculate how far apart are the birds.




1
Expert's answer
2021-11-29T15:13:53-0500

"\theta=\tan \theta=\frac hL=\frac{1.22\lambda}D,"

"h=\frac{1.22\lambda L}D=1~cm."